ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

14. Mathematical Aspects of Computer Science

1991 MS Classification: 90c60 Gärtner, Bernd, ETH Zürich, Switzerland Abstract Objective Functions - From Geometry to Combinatorics in Linear Programming Over the last six years, generalizations of linear programming have successfully been used in the analysis of randomized simplex algorithms. We survey some of the recent results and address the role of combinatorics in linear programming with respect to the simplex method.